OpenStudy (anonymous):

If the Wronskian of f and g is tcost-sint, and if u=f+3g, v=f-g, find the Wronskian of u and v. ( I know that (f+3g)(f-g)'-(f+3g)'(f-g) but what do I do after that? Please help me.)

OpenStudy (loser66):

=-4(tcost-sint).just calculate as usual.

OpenStudy (anonymous):

But how?

OpenStudy (loser66):

\[\left[\begin{matrix}u &v \\ u' & v'\end{matrix}\right]=\left[\begin{matrix}(f+3g) & (f-g) \\( f'+3g') & (f'-g')\end{matrix}\right]\]

OpenStudy (loser66):

take determinant of that matrix, you will see it cancel out and just -4(Wronskian of f and g)

OpenStudy (anonymous):

But isn't it there's a way to do it without doing the matrix? Because I did (f+3g)(1-1)-(1+3)(f-g)=-4(f-g), is this right?

OpenStudy (loser66):

you are kicked out of the net? I guess!! ok, I calculate it for you

OpenStudy (loser66):

\[(f+3g)(f'-g')-(f'+3g')(f-g) \\= (ff'-fg'+3gf'-3gg')-(ff'-gf'+3fg'-3gg')\\=ff'-fg'+3gf'-3gg'-ff'+f'g-3fg'+3gg'\] \[=4gf'-4fg'\\=-4(fg'-gf')\\=-4Wronskian(f,g)\\=-4(tcost-sint)\]

OpenStudy (anonymous):

Wait a minute.

OpenStudy (anonymous):

Thank you, you were right.

OpenStudy (loser66):

heehe, it's not hard for you, too. just stuck with relationship between f and g and get nowhere to go.